OpenStudy (lisa123):

Find two positive real numbers whose product is a maximum. (Enter your answers as a comma-separated list

OpenStudy (lisa123):

the sum of the first and twice the second is 20

zepdrix (zepdrix):

Hey Lisa! :) What type of class is this for? Can we use Calculus Optimization methods?

OpenStudy (lisa123):

Hey! and no this is precalc with algebra

zepdrix (zepdrix):

Ooo that's no fun >.< I'm trying remember how to do these optimize probems without calculus... Hmmm

OpenStudy (lisa123):

There is one problem I did where it said the sum is equal to 130 So I set up an equation of x+y=130 and then I divided both sides by 2 and got 65,65 as my answer

OpenStudy (lisa123):

I doubt my reasoning is right but I got the correct answer

jimthompson5910 (jim_thompson5910):

first isolate y in `x+y=130` to get `y = -x+130`

jimthompson5910 (jim_thompson5910):

so `x*y` turns into `x*(-x+130) = -x^2 + 130x`

jimthompson5910 (jim_thompson5910):

finding the vertex of `f(x) = -x^2 + 130x` will help you find the max product and the two values to get the max product

OpenStudy (lisa123):

Oh wow okay that makes sense how 65 is the answer

jimthompson5910 (jim_thompson5910):

`-x^2 + 130x` is in the form `ax^2 + bx + c ` a = -1 b = 130 compute `h = -b/(2a)` to get `h = 65`. Then plug this into `-x^2 + 130x` to find the product

OpenStudy (lisa123):

Okay I see. How would I set up the equation for the first problem?

jimthompson5910 (jim_thompson5910):

This one? `the sum of the first and twice the second is 20`

OpenStudy (lisa123):

Yes

OpenStudy (lisa123):

would it be x+2y=20?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

then you can isolate x to get x = -2y+20 so x*y = (-2y+20)*y = -2y^2 + 20y then treat y as x (they're both variables, so why not) find the vertex of -2x^2 + 20x

OpenStudy (lisa123):

= 5

jimthompson5910 (jim_thompson5910):

so y is really 5 since I replaced y with x y = 5 x = -2y+20 x = -2*5+20 x = -10+20 x = 10

OpenStudy (lisa123):

ohhh wow I understand. Thank you so much for helping me @jim_thompson5910 !!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (lisa123):

Do you mind helping me with another question?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (lisa123):

I have to write the quadratic function x^2-10x in standard form

jimthompson5910 (jim_thompson5910):

do you mean in vertex form?

OpenStudy (lisa123):

no the directions say standard.....I think you have to convert it into vertex to get the standard

jimthompson5910 (jim_thompson5910):

does your book say that standard form is ax^2 + bx + c? or y - k = a(x-h)^2 ?

OpenStudy (lisa123):

first one

jimthompson5910 (jim_thompson5910):

well x^2-10x is already in standard form look at how it matches up to ax^2 + bx + c in this case a = 1 b = -10 c = 0

OpenStudy (lisa123):

no...when I plug it in it tells me its wrong so I guess the second one

jimthompson5910 (jim_thompson5910):

ok so you need to find the vertex

jimthompson5910 (jim_thompson5910):

you first need to compute h = -b/(2a)

OpenStudy (lisa123):

h=5

jimthompson5910 (jim_thompson5910):

yes, now plug that back into the given function to find y

OpenStudy (lisa123):

25-50= -25 y=-25

jimthompson5910 (jim_thompson5910):

yep the vertex is (5,-25)

jimthompson5910 (jim_thompson5910):

so (h,k) = (5,-25) h = 5 k = -25

jimthompson5910 (jim_thompson5910):

a = 1 h = h k = -25 plug that into `y - k = a(x-h)^2`

OpenStudy (lisa123):

-50=(x-5)^2...?

jimthompson5910 (jim_thompson5910):

you should get `y+25 = (x-5)^2`

OpenStudy (lisa123):

okay then what?

jimthompson5910 (jim_thompson5910):

that's the answer in `y-k = a(x-h)^2` form